# Extrinsic contribution to bosonic thermal Hall transport

### v1

@@ VI Conclusion
We have shown that the side-jump contribution gives a sizable non-perturbative contribution to the bosonic THE. As a side product, we also find a generalized formula of the side-jump contribution to the electronic AHE from generic types of impurities (e.g., any local deformation of the Hamiltonian), see Eq. IV.5. This contribution is an order one correction to the intrinsic part of the Hall response, that is independent of the density of impurities only if the carriers’ mean free path is governed by disorder scattering. For sufficiently strong interactions, such that the mean free path is independent of impurity density, the side-jump contribution becomes perturbative ($\propto n_{\rm imp}$) and might be negligible – the price to pay is that interactions, in turn, will play a role in the Hall conductivity, e.g. via many-body skew-scattering mechanisms. In particular, should there be a near-“universal” mechanism for THE in trivial non-magnetic insulators, interactions should be involved to some extent.
Beyond thermal Hall transport, our gauge invariant formulation of a kinetic equation for bosons including disorder scattering promises many useful applications. For instance it naturally allows for the calculation of other transport coefficients, such as the spin Nernst effect. It also points to several concrete directions for future research: (1) The kinetic formalism should be extended for calculating nonlinear response functions – for a recent treatment of the clean case see Park et al. (2025). (2) It can also be generalized to include time-dependent (random) terms in the Hamiltonian, allowing in particular to study in a minimal version the dissipative quantum dynamics of systems with nontrivial band geometry. (3) Our results indicate that quantitative predictions for the THE in the intrinsic regime will require knowledge of the operatorial form of disorder. Thus, it will be useful to interface our theory with quantitative ab-initio efforts to obtain not only the clean features (energies, wavefunctions) of the phonons or magnons, but also input for the effective disorder Hamiltonian. (4) A challenging next step will be the inclusion of interaction effects between bosons, which will not only provide yet another contribution to the thermal Hall effect from skew-scattering, but will also alter the side-jump contribution from impurities, and may even generate a side-jump-like contribution of their own by mixing many-body dynamics with band geometrical properties.
@@ Acknowledgements
L.M. thanks Elio König for a useful discussion. L.M. thanks Lucile Savary and Leon Balents for helpful comments and a previous collaboration that led to this work. We thank Peng Rao for useful discussion and acknowledge support from the Imperial-TUM flagship partnership, from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy–EXC– 2111–390814868, DFG grants No. KN1254/1-2, KN1254/2-1, and TRR 360 - 492547816, as well as the Munich Quantum Valley, which is supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus.
@@ Appendix A Derivation of the kinetic equation: calculatory details
×$\qquad\qquad\times\left[\mathsf{S}^{-1}(p+k){\mathsf{W}}(p+k/2)\mathsf{S}(p)\right]^{s}_{n,n^{\prime}}\;\left[\mathsf{S}^{-1}(p^{\prime}){\mathsf{W}}(p^{\prime}+k/2)\mathsf{S}(p^{\prime}+k)\right]^{s^{\prime}}_{n^{\prime},n}.$
@@
×$\times\Big{[}\langle(\mathsf{V}_{\rm d})_{n,n^{\prime}}(x,u)(\mathsf{V}_{\rm d})_{n^{\prime},n}(\xi,\zeta)\rangle\delta_{(\zeta^{\prime}-\xi^{\prime})}\delta_{(v-y)}\delta_{l,n^{\prime}}\delta_{l^{\prime},n}\delta_{m,n}-\langle(\mathsf{V}_{\rm d})_{n,n^{\prime}}(\xi,\zeta)(\mathsf{V}_{\rm d}^{\dagger})_{n^{\prime},n}(v,y)\rangle\delta_{(x-u)}\delta_{(\zeta^{\prime}-\xi^{\prime})}\delta_{l,n}\delta_{l^{\prime},n^{\prime}}\delta_{m,n^{\prime}}$
@@
×$\times\left[\mathsf{S}^{-1}(p+K/2)\mathsf{W}(\tfrac{p+p^{\prime}}{2}+K/2)\mathsf{S}(p^{\prime}+K/2)\right]_{n,n^{\prime}}\left[\mathsf{S}^{-1}(p^{\prime}+K/2-\Pi)\mathsf{W}(\tfrac{p+p^{\prime}}{2}+K/2-\Pi)\mathsf{S}(p+K/2-\Pi)\right]_{n^{\prime},n}$
@@
×$\times\left[\mathsf{S}^{-1}(p+K/2)\mathsf{W}(\tfrac{p+p^{\prime}}{2}+K/2)\mathsf{S}(p^{\prime}+K/2)\right]_{n,n^{\prime}}\left[\mathsf{S}^{-1}(p-K/2-\Pi^{\prime})\mathsf{W}(\tfrac{p+p^{\prime}}{2}-K/2-\Pi^{\prime})\mathsf{S}(p^{\prime}-K/2-\Pi^{\prime})\right]^{\dagger}_{n^{\prime},n}$
@@
×$\times\left[\mathsf{S}^{-1}(p+K/2)\mathsf{W}(\tfrac{p+p^{\prime}}{2}+K/2)\mathsf{S}(p^{\prime}+K/2)\right]_{n,n^{\prime}}\left[\mathsf{S}^{-1}(p-K/2-\Pi)\mathsf{W}(\tfrac{p+p^{\prime}}{2}-K/2-\Pi)\mathsf{S}(p^{\prime}-K/2-\Pi)\right]^{\dagger}_{n^{\prime},n}$
@@
×$\times\left[\mathsf{S}^{-1}(p^{\prime}-K/2)\mathsf{W}(\tfrac{p+p^{\prime}}{2}-K/2)\mathsf{S}(p-K/2)\right]^{\dagger}_{n,n^{\prime}}\left[\mathsf{S}^{-1}(p-K/2-\Pi^{\prime})\mathsf{W}(\tfrac{p+p^{\prime}}{2}-K/2-\Pi^{\prime})\mathsf{S}(p^{\prime}-K/2-\Pi^{\prime})\right]^{\dagger}_{n^{\prime},n}\Bigg{\}}.$
We now perform the Wigner transform with respect to $(x,y)$ on both sides, by writing $X=\frac{x+y}{2}$ and performing the Fourier transform with respect to $x-y$. We further perform one momentum integral, resolve one momentum conservation delta-function, identify the matrix elements $\mathsf{M}_{n,n^{\prime}}(p,p^{\prime})$, perform momentum shifts by amounts ±$\pm\Pi,\pm\Pi^{\prime}$ (different for each term in the sum) and neglect such $K,\Pi,\Pi^{\prime}$ dependence in the second (i.e. fast) variable of $\breve{\mathsf{F}}_{\rm d},\breve{\mathsf{K}}_{\rm d}$ as explained in the main text, Sec. III.1.
@@
×$\times\sum_{n^{\prime}}\Big{\{}e^{-\frac{i}{\hbar}(t-\tau)((\breve{\mathsf{K}}_{\rm d})_{n^{\prime},n^{\prime}}(\Pi,p^{\prime})-(\breve{\mathsf{K}}_{\rm d})_{n,n}(\Pi^{\prime},p))}$
@@
×$\qquad\qquad\qquad\times\mathsf{M}_{n,n^{\prime}}(p+\tfrac{K}{2}+\tfrac{\Pi^{\prime}}{2},p^{\prime}+\tfrac{K}{2}+\tfrac{\Pi^{\prime}}{2})\;\mathsf{M}_{n,n^{\prime}}(p-\tfrac{K}{2}-\tfrac{\Pi^{\prime}}{2},p^{\prime}-\tfrac{K}{2}-\tfrac{\Pi^{\prime}}{2})^{*}\;(\breve{\mathsf{F}}_{\rm d})_{n^{\prime},n^{\prime}}(K,p^{\prime},\tau)\Big{\}}_{\begin{subarray}{c}p\rightarrow p-(\Pi-\Pi^{\prime})/2,\\ p^{\prime}\rightarrow p^{\prime}-(\Pi-\Pi^{\prime})/2\end{subarray}}.$
@@
×$\qquad\qquad\qquad\qquad\qquad\qquad\times\delta\left((\mathsf{K}_{\rm d})_{n^{\prime},n^{\prime}}(X+\delta{\mathsf{r}}_{n,n^{\prime}}^{-}(p,p^{\prime})-\delta{\mathsf{r}}_{n,n^{\prime}}^{+}(p,p^{\prime}),p^{\prime})-(\mathsf{K}_{\rm d})_{n,n}(X+\delta{\mathsf{r}}_{n,n^{\prime}}^{+}(p,p^{\prime}),p)\right)$
@@
×$\qquad\qquad\qquad\qquad\qquad\times\delta\left((\mathsf{K}_{\rm d})_{n^{\prime},n^{\prime}}(X+\delta{\mathsf{r}}_{n,n^{\prime}}^{+}(p,p^{\prime})+\delta{\mathsf{r}}_{n,n^{\prime}}^{-}(p,p^{\prime}),p^{\prime})-(\mathsf{K}_{\rm d})_{n,n}(X,p)\right)\bigg{]},$
@@ B.1 Obtaining the disorder-induced curvature
×$\times\bigg{(}1+(p^{\prime}_{\nu}-p_{\nu})\frac{\left[\mathsf{S}^{-1}(p)\mathsf{W}(p)\partial_{p_{\nu}}\mathsf{S}(p)\right]_{n,n^{\prime}}}{\left[\mathsf{S}^{-1}(p)\mathsf{W}(p)\mathsf{S}(p)\right]_{n,n^{\prime}}}+\tfrac{1}{2}(p^{\prime}_{\nu}-p_{\nu})\frac{\left[\mathsf{S}^{-1}(p)\partial_{p_{\nu}}\mathsf{W}(p)\mathsf{S}(p)\right]_{n,n^{\prime}}}{\left[\mathsf{S}^{-1}(p)\mathsf{W}(p)\mathsf{S}(p)\right]_{n,n^{\prime}}}\bigg{)},$ (64)
@@
Here is is easy to check gauge invariance under $-ϑ\mathsf{S}\rightarrow\mathsf{S}e^{i\stheta}$ of $\delta\mathsf{X}^{\mu}_{n,n^{\prime}}(p,p^{\prime})$ defined in Eq.(26). Indeed, the gauge transformation rules are $\mathsf{A}^{\mu}_{n}(p)\rightarrow\mathsf{A}^{\mu}_{n}(p)+\partial_{p_{\mu}}\theta_{n}(p)$ and $\delta x^{\mu}_{n,n^{\prime}}(p)\rightarrow\delta x^{\mu}_{n,n^{\prime}}(p)-p_{\nu}\partial^{2}_{\mu\nu}\theta_{n^{\prime}}$ (note the second term in $\delta x^{\mu}_{n,n^{\prime}}(p)$ is gauge-invariant by itself), and
@@ C.1 Neglecting $\delta\mathsf{r}^{+}$:
We first note that $\delta\mathsf{r}^{+}$ is gauge-invariant, therefore taking $\delta\mathsf{r}^{+}\rightarrow 0$ is a well-defined approximation. Note that in the scalar fermionic case, $\delta\mathsf{r}^{+}=0$ is rigorous. We now argue why it can also be neglected in the cases we are considering in this paper.
@@ C.2 Neglecting interband $(n\neq n^{\prime})$ terms:
at least for $n=n^{\prime}$ (see App.C.2) and for all $p$. We note that because matrix elements $\mathsf{M}_{n,n^{\prime}}(p,p^{\prime})$ appear pairwise in the collision integral, the condition “$\in\mathbb{R}$” (instead of “$>0$”) is sufficient.
In the particular case $\mathsf{W}=\mathsf{K}$, the validity of Eq.(73) follows directly from Eq.(71). More generally, all the cases we consider for applications (see Appendix G) satisfy this constraint, although this is not a priori a general truth. This justifies the replacement here.
@@ D.2 Semiclassical expansion
×$\times\delta\mathsf{X}_{n,n^{\prime}}(p,p^{\prime})\;(\underline{\mathsf{F}}_{\rm d})_{n^{\prime},n^{\prime}}(X+\delta\mathsf{X}_{n,n^{\prime}}(p,p^{\prime}),p^{\prime},t)\;\left|\mathsf{M}_{n,n^{\prime}}(p,p^{\prime})\right|^{2}\;\tilde{w}(p-p^{\prime}),$ (80)
@@
×$\times\bigg{[}\big{(}X-X\big{)}\times(\underline{\mathsf{F}}_{\rm d})_{n,n}(X,p,t)\;\left|\mathsf{M}_{n,n^{\prime}}(p,p^{\prime})\right|\,\left|\mathsf{M}_{n^{\prime},n}(p^{\prime},p)\right|$

### v2

@@ VI Conclusion
We have derived systematically the side-jump contribution to the bosonic THE, Eq.(2). As a side product, we also found a generalized formula of the side-jump contribution to the electronic AHE from generic types of impurities (i.e., any local deformation of the Hamiltonian), see Eq. IV.5. This contribution is a sizeable correction to the intrinsic part of the Hall response, that is non-perturbative i.e. independent of the density of impurities (in the absence of interactions). Thus, the prevailing interpretation of thermal Hall experiments with Eq.(1) alone is always incomplete, and we have shown quantitatively (both formally and in a realistic spin model for $\alpha\text{-RuCl}_{3}$) that extrinsic contributions can be of similar magnitude. This is especially so in bosonic systems, where so far Eq.(1) has been most commonly used alone. Therefore, we argue that any reasonable estimate of the THE in the absence of interactions should involve both Eq.(1) and Eq.(2) together – the latter can be evaluated straightforwardly, for instance using the procedure detailed in App. G.
This independence of Eq.(2) on the impurity density holds only if the carriers’ mean free path is governed by disorder scattering. For sufficiently strong interactions, such that the mean free path is independent of impurity density, the side-jump contribution becomes perturbative ($\propto n_{\rm imp}$) and might be negligible – the price to pay is that interactions, in turn, will play a role in the Hall conductivity, e.g. via many-body skew-scattering mechanisms. Ultimately, for bosonic systems there is no reason to believe that Eq.(1) alone should ever provide a good approximation of the thermal Hall conductivity in any experiment. In particular, should there be a near-“universal” mechanism for THE in trivial non-magnetic insulators Behnia (2025), interactions should be involved to some extent.
Beyond thermal Hall transport, our gauge invariant formulation of a kinetic equation for bosons including disorder scattering promises many useful applications. For instance it naturally allows for the calculation of other transport coefficients, such as the spin Nernst effect. It also points to several concrete directions for future research: (1) The kinetic formalism should be extended for calculating nonlinear response functions – for a recent treatment of the clean case see Park et al. (2025). (2) It can also be generalized to include time-dependent (random) terms in the Hamiltonian, allowing in particular to study in a minimal version the dissipative quantum dynamics of systems with nontrivial band geometry. (3) Our results indicate that quantitative predictions for the THE in the intrinsic regime will require knowledge of the operatorial form of disorder. Thus, it will be useful to interface our theory with quantitative ab-initio efforts to obtain not only the clean features (energies, wavefunctions) of the phonons or magnons Dhakal et al. (2024), but also input for the effective disorder Hamiltonian, as we describe in Sec.III.5. (4) A challenging next step will be the inclusion of interaction effects between bosons, which will not only provide yet another contribution to the thermal Hall effect from skew-scattering, but will also alter the side-jump contribution from impurities, and may even generate a side-jump-like contribution of their own by mixing many-body dynamics with band geometrical properties.
@@ Acknowledgements
L.M. thanks Elio König and Joji Nasu for a useful discussion. L.M. thanks Lucile Savary and Leon Balents for helpful comments and a previous collaboration that led to this work. We thank Peng Rao for useful discussion and acknowledge support from the Imperial-TUM flagship partnership, from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy–EXC– 2111–390814868, DFG grants No. KN1254/1-2, KN1254/2-1, and TRR 360 - 492547816, as well as the Munich Quantum Valley, which is supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus.
@@ Appendix A Derivation of the kinetic equation: calculatory details
$\qquad\qquad\times\left[\mathsf{S}^{-1}(p+k){\mathsf{W}}(p+k/2)\mathsf{S}(p)\right]^{s}_{n,n^{\prime}}\;\left[\mathsf{S}^{-1}(p^{\prime}){\mathsf{W}}(p^{\prime}+k/2)\mathsf{S}(p^{\prime}+k)\right]^{s^{\prime}}_{n^{\prime},n}.$
@@
$\times\Big[\langle(\mathsf{V}_{\rm d})_{n,n^{\prime}}(x,u)(\mathsf{V}_{\rm d})_{n^{\prime},n}(\xi,\zeta)\rangle\delta_{(\zeta^{\prime}-\xi^{\prime})}\delta_{(v-y)}\delta_{l,n^{\prime}}\delta_{l^{\prime},n}\delta_{m,n}-\langle(\mathsf{V}_{\rm d})_{n,n^{\prime}}(\xi,\zeta)(\mathsf{V}_{\rm d}^{\dagger})_{n^{\prime},n}(v,y)\rangle\delta_{(x-u)}\delta_{(\zeta^{\prime}-\xi^{\prime})}\delta_{l,n}\delta_{l^{\prime},n^{\prime}}\delta_{m,n^{\prime}}$
@@
$\times\left[\mathsf{S}^{-1}(p+K/2)\mathsf{W}(\tfrac{p+p^{\prime}}{2}+K/2)\mathsf{S}(p^{\prime}+K/2)\right]_{n,n^{\prime}}\left[\mathsf{S}^{-1}(p^{\prime}+K/2-\Pi)\mathsf{W}(\tfrac{p+p^{\prime}}{2}+K/2-\Pi)\mathsf{S}(p+K/2-\Pi)\right]_{n^{\prime},n}$
@@
$\times\left[\mathsf{S}^{-1}(p+K/2)\mathsf{W}(\tfrac{p+p^{\prime}}{2}+K/2)\mathsf{S}(p^{\prime}+K/2)\right]_{n,n^{\prime}}\left[\mathsf{S}^{-1}(p-K/2-\Pi^{\prime})\mathsf{W}(\tfrac{p+p^{\prime}}{2}-K/2-\Pi^{\prime})\mathsf{S}(p^{\prime}-K/2-\Pi^{\prime})\right]^{\dagger}_{n^{\prime},n}$
@@
$\times\left[\mathsf{S}^{-1}(p+K/2)\mathsf{W}(\tfrac{p+p^{\prime}}{2}+K/2)\mathsf{S}(p^{\prime}+K/2)\right]_{n,n^{\prime}}\left[\mathsf{S}^{-1}(p-K/2-\Pi)\mathsf{W}(\tfrac{p+p^{\prime}}{2}-K/2-\Pi)\mathsf{S}(p^{\prime}-K/2-\Pi)\right]^{\dagger}_{n^{\prime},n}$
@@
$\times\left[\mathsf{S}^{-1}(p^{\prime}-K/2)\mathsf{W}(\tfrac{p+p^{\prime}}{2}-K/2)\mathsf{S}(p-K/2)\right]^{\dagger}_{n,n^{\prime}}\left[\mathsf{S}^{-1}(p-K/2-\Pi^{\prime})\mathsf{W}(\tfrac{p+p^{\prime}}{2}-K/2-\Pi^{\prime})\mathsf{S}(p^{\prime}-K/2-\Pi^{\prime})\right]^{\dagger}_{n^{\prime},n}\Bigg\}.$
We now perform the Wigner transform with respect to $(x,y)$ on both sides, by writing $X=\frac{x+y}{2}$ and performing the Fourier transform with respect to xyx-y. We further perform one momentum integral, resolve one momentum conservation delta-function, identify the matrix elements $\mathsf{M}_{n,n^{\prime}}(p,p^{\prime})$, perform momentum shifts by amounts $\pm\Pi,\pm\Pi^{\prime}$ (different for each term in the sum) and neglect such $K,\Pi,\Pi^{\prime}$ dependence in the second (i.e. fast) variable of $\breve{\mathsf{F}}_{\rm d},\breve{\mathsf{K}}_{\rm d}$ as explained in the main text, Sec. III.1.
@@
$\times\sum_{n^{\prime}}\Big\{e^{-\frac{i}{\hbar}(t-\tau)((\breve{\mathsf{K}}_{\rm d})_{n^{\prime},n^{\prime}}(\Pi,p^{\prime})-(\breve{\mathsf{K}}_{\rm d})_{n,n}(\Pi^{\prime},p))}$
@@
$\qquad\qquad\qquad\times\mathsf{M}_{n,n^{\prime}}(p+\tfrac{K}{2}+\tfrac{\Pi^{\prime}}{2},p^{\prime}+\tfrac{K}{2}+\tfrac{\Pi^{\prime}}{2})\;\mathsf{M}_{n,n^{\prime}}(p-\tfrac{K}{2}-\tfrac{\Pi^{\prime}}{2},p^{\prime}-\tfrac{K}{2}-\tfrac{\Pi^{\prime}}{2})^{*}\;(\breve{\mathsf{F}}_{\rm d})_{n^{\prime},n^{\prime}}(K,p^{\prime},\tau)\Big\}_{\begin{subarray}{c}p\rightarrow p-(\Pi-\Pi^{\prime})/2,\\ p^{\prime}\rightarrow p^{\prime}-(\Pi-\Pi^{\prime})/2\end{subarray}}.$
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\times\delta\left((\mathsf{K}_{\rm d})_{n^{\prime},n^{\prime}}(X+\delta{\mathsf{r}}_{n,n^{\prime}}^{-}(p,p^{\prime})-\delta{\mathsf{r}}_{n,n^{\prime}}^{+}(p,p^{\prime}),p^{\prime})-(\mathsf{K}_{\rm d})_{n,n}(X+\delta{\mathsf{r}}_{n,n^{\prime}}^{+}(p,p^{\prime}),p)\right)$
@@
$\qquad\qquad\qquad\qquad\qquad\times\delta\left((\mathsf{K}_{\rm d})_{n^{\prime},n^{\prime}}(X+\delta{\mathsf{r}}_{n,n^{\prime}}^{+}(p,p^{\prime})+\delta{\mathsf{r}}_{n,n^{\prime}}^{-}(p,p^{\prime}),p^{\prime})-(\mathsf{K}_{\rm d})_{n,n}(X,p)\right)\bigg],$
@@ B.1 Obtaining the disorder-induced curvature
$\times\bigg(1+(p^{\prime}_{\nu}-p_{\nu})\frac{\left[\mathsf{S}^{-1}(p)\mathsf{W}(p)\partial_{p_{\nu}}\mathsf{S}(p)\right]_{n,n^{\prime}}}{\left[\mathsf{S}^{-1}(p)\mathsf{W}(p)\mathsf{S}(p)\right]_{n,n^{\prime}}}+\tfrac{1}{2}(p^{\prime}_{\nu}-p_{\nu})\frac{\left[\mathsf{S}^{-1}(p)\partial_{p_{\nu}}\mathsf{W}(p)\mathsf{S}(p)\right]_{n,n^{\prime}}}{\left[\mathsf{S}^{-1}(p)\mathsf{W}(p)\mathsf{S}(p)\right]_{n,n^{\prime}}}\bigg),$ (75)
@@
[…]
@@ C.1 Neglecting $\delta\mathsf{r}^{+}$:
We first note that $\delta\mathsf{r}^{+}$ is gauge-invariant, therefore taking $\delta\mathsf{r}^{+}\rightarrow 0$ is a well-defined approximation. Note that in the case of fermions in a random scalar potential, $\delta\mathsf{r}^{+}=0$ is rigorous. We now argue why it can also be neglected in the cases we are considering in this paper.
@@ C.2 Neglecting interband $(n\neq n^{\prime})$ terms:
at least for $n=n^{\prime}$ (see App.C.2) and for all $p$. We note that because matrix elements $\mathsf{M}_{n,n^{\prime}}(p,p^{\prime})$ appear pairwise in the collision integral, the condition $\in\mathbb{R}$” (instead of “>0>0”) is sufficient.
In the particular case $\mathsf{W}=\mathsf{K}$, the validity of Eq.(84) follows directly from Eq.(82). More generally, all the cases we consider for applications (see Appendix H) satisfy this constraint, although this is not a priori a general truth. This justifies the replacement here.
@@ D.2 Semiclassical expansion
$\times\delta\mathsf{X}_{n,n^{\prime}}(p,p^{\prime})\;(\underline{\mathsf{F}}_{\rm d})_{n^{\prime},n^{\prime}}(X+\delta\mathsf{X}_{n,n^{\prime}}(p,p^{\prime}),p^{\prime},t)\;\left|\mathsf{M}_{n,n^{\prime}}(p,p^{\prime})\right|^{2}\;\tilde{w}(p-p^{\prime}),$ (91)
@@
$\times\bigg[\big(X-X\big)\times(\underline{\mathsf{F}}_{\rm d})_{n,n}(X,p,t)\;\left|\mathsf{M}_{n,n^{\prime}}(p,p^{\prime})\right|\,\left|\mathsf{M}_{n^{\prime},n}(p^{\prime},p)\right|$